I had a question with 6) B( 20, .7) P(10 < x < 16) =P(11< or equal x < or equal 15) = P (x less or equal 15)-P(x less or equal 10) .762-.048= .714 .714 is the answer anyone help
Answer #1

B(20, 0.7)

n = 20

p = 0.7

P(10 <x<16) = P(X<16) - P(X<=10)

= P(X<=15) - P(X<=10)

Using binomial function in excel

= BINOM.DIST(15, 20, 0.7, 1) - BINOM.DIST(10, 20, 0.7, 1)

= 0.7625 - 0.0480

= 0.7145

Yes.